OpenStudy (anonymous):

Which system of linear equations has no solution? A. 2x − y = −1 2x + y = 2 B. 4x − y = 3 3x − 2y = 3 C. −2x + 4y = −16 4x − 2y = −1 D. 2x − y = −1 −2x + y = −1 D!?!?!?

OpenStudy (anonymous):

@jim_thompson5910 @hartnn @phi @zepdrix @thomaster @tHe_FiZiCx99 @TheRealMeeeee @mathslover @iambatman @satellite73 @Preetha

mathslover (mathslover):

Well done!

OpenStudy (anonymous):

What about this one? :) Which system of linear equations shares the same line? A? @mathslover A. 6x − 3y = 15 8x − 4y = 20 B. 6x − 2y = 2 x 1/3 + y = 1/3 C. 3x − 3y = 15 9x − 6y = 6 D. −4x + 4y = −4 2x + 2y = 2

mathslover (mathslover):

Good Work!

OpenStudy (anonymous):

Using linear combinations, what is the solution(s) to the following system of linear equations, if there is one? −3x + 4y = −1 4x − 4y = −4 A. No solution B. (−3, −2 1/2) C. (−5, −4) D. (−4, −5) @mathslover

OpenStudy (anonymous):

Can someone erxplain it to me? @mathslover @tHe_FiZiCx99 @Hero @phi @Preetha

mathslover (mathslover):

What is special about this system of linear equations? The given system of equations, has something to cancel out. try to add the 2 equations, see what you get!

OpenStudy (the_fizicx99):

Linear combinations, hmm. It's been a while since I've done them but it's basically using one of the terms to multiply the other equation so they both have a common term to cancel out. You can already cancel out one term by adding them.

OpenStudy (anonymous):

what do i do with the x and y?

OpenStudy (the_fizicx99):

-3x + 4x = ? 4y + (-4y) = ?

OpenStudy (the_fizicx99):

Don't forget the constants either.

OpenStudy (anonymous):

They already answered that. -1 and -4. @tHe_FiZiCx99

OpenStudy (the_fizicx99):

o_o 4y + (-4y) = -4 ? >.> They cancel out, you're left with x = -5 Use "x" to solve for "y" by subbing in -5 for "x" You can chose to sub it into either equation.

hero (hero):

@CrazyCountryGirl, what steps did you take to solve the first two systems? Can you explain your approaches here?

OpenStudy (anonymous):

What? Oh! Those aren't mine. They're my friends. I let her borrow my account for some practice home work her teacher gave her to do over the summer. Os wouldn't let me close the question. I posted the third. Sorry for the confusion. :)

hero (hero):

FYI: Never allow anyone else to use your account. Tell your friend to create her own account. Such statements discourages people from wanting to help you.

OpenStudy (anonymous):

@Hero It doesn't say in the rules that we can't, that I know of. o_o @tHe_FiZiCx99 Can you write that as an equation? Like replace the x with -5 and stuff? :) It might be easier for me to work out. :):):)

hero (hero):

There's also technically no rule for trolling but users get suspended for trolling everyday.

OpenStudy (anonymous):

oh... o_o

OpenStudy (the_fizicx99):

Ehm, \(\ \sf x = \color{blue}{ -5 } \) \(\ \sf −3(\color{blue}{-5}) + 4y = −1 \) \(\ \sf 15 + 4\color{red}{y} = -1 \) Solve for \(\ \sf "\color{red}{y}" \)

hero (hero):

By the way, just because something isn't on the "rules" list, that doesn't mean it is okay to do it. Use your WISE judgment to make the right decisions at any given moment.

hero (hero):

In most cases, common sense applies.

OpenStudy (anonymous):

@tHe_FiZiCx99 -4?

OpenStudy (anonymous):

c or d?

OpenStudy (anonymous):

C? I positive it's c!

OpenStudy (the_fizicx99):

Yes. Put it in an ordered pair. \(\ \sf (\color{blue}{x},\color{red}{y}) \)

OpenStudy (anonymous):

cause x before y, and x=-5